12 GeV Upgrade
Both of the coldboxes for the new Central Helium Liquefier (CHL-2) have arrived and have been put into final positions. The C100-1 and C100-2 cryomodules continue to be used for beam operations at voltages up to 91 MV; refinement of the control of the radiofrequency fields continues. Helium headers have been attached for the eighth and ninth C100 cryomodules. Installation of all klystrons for the new RF zones in the South Linac is complete. Adjustments to the FY12 and FY13 accelerator work plans, in response to the reduction in FY12 funding relative to the baseline profile, are nearing completion.
Last week, the g2p experiment in Hall A started production running and finished the calibrations and commissioning. Unfortunately, the septum magnet had an electrical failure, causing significant damage to the magnet on Saturday. The extent of this damage and the path forward are being evaluated at this time.
In Hall C, the Q-weak experiment took production data on hydrogen and some background data with an aluminum target. The experiment started using the halo target, a target with a hole for the beam to pass through, to warn of poor beam conditions.
Minority Undergraduate Research Assistantship
Last week began with a spot move: the spot on the cathode where the lasers hit to produce electrons was moved to an area on the cathode that produces electrons more efficiently. On Tuesday, a faulty valve interrupted cooling water flow to some injector components, requiring an access downstairs to the injector to investigate and repair. The injector experiment, PEPPo, performed tests using the injector beam while software and firmware upgrades were installed and worked on in the new C100 accelerating cavities in the South Linac. After beam studies wrapped up and a few repairs were made, beam was restored to Halls A and C, while Hall B continued target work. At the end of the week, zone 1L02 had a problem maintaining its liquid helium level due poor insulating vacuum. A supplemental vacuum pump was installed, which took a good chunk out of Friday's beam time. The weekend was quiet, except for Hall A issues with a septum magnet.
Operations of the FEL continued to optimize the machine for optics measurements, which are beginning this week. During the early shift, FEL staffers continued to assemble and align components of the ultraviolet wiggler chamber. The Office of Naval Research announced the cancellation of the Boeing FEL-Innovative Naval Prototypes program effort to happen after this week's conceptual design report.
Center for Theoretical and Computational Physics
Last week, a small group of experimentalists and theorists came together for a four-day workshop on Twin Approaches to Confinement Physics to consider how the fact that quarks are confined inside hadrons (and have never been found alone) determines the detailed properties of hadrons: their spectrum, structure and dynamics. The workshop discussed how experiments at JLab can probe the very mechanism of confinement and help us understand how QCD really works.
The Electrical Engineering Support Radiofrequency group completed testing of zone SL23, the third new 12 GeV RF system, at the end of February. Testing on SL22, the fourth new RF system, has begun. The RF Group by now has received nearly 70 percent of the the new high-power RF tubes (klystrons) needed for the 12 GeV upgrade. The EES DC Power, Instrumentation and Control, and Safety System groups have been busy making preparations for the upcoming long shutdown that begins in May. The EES fabrication shop has completed the assembly of all the remaining RF control crates needed for 12 GeV. They are currently assembling beam position monitor electronics, Personnel Safety System run/safe boxes and machine fast shutdown protection hardware for installation later this year. The EES electricians have worked on AC power infrastructure installation and upgrades around the accelerator in preparation for all the new equipment planned. All groups continue to provide accelerator operations support for the current physics run as well.
Environment, Safety, Health and Quality
Safety First When Walking/Jogging Onsite
Last Chance to Vote for Your Favorite JLab 2012 T-Shirt Design!
Volunteers Needed for May 19 JLab Open House
National Rent-A-Car Promotion Onsite Thursday
Seminar on Women in Physics Offered April 5
12th FLUKA Course Offered at JLab April 30 - May 4
Tracking Thomas at JLab
Congratulations this week go to David Lawrence, Samika Hawkins and Michelle Shinn, who were the first to correctly identify the March 14 location. Honorable mentions go to Michael Barnes, Michael Haddox-Schatz, Subashini De Silva, Doug Higinbotham and Dick Owen. Check out the Tracking Thomas webpage for a better view of his last location and this week's new mystery photo.
JLab Calendar of Events
March 26: Safety Shoe Vendor onsite